If two fair dice are rolled, what is the conditional probability that the rst one lands on an even number given that the second one lands on a number less than or equal to four? Compute the conditional probability that the second one lands on a number less than four given that the rst one lands on an even number.
Answer #1

Since the two events that the first die lands on an even number and the second die lands on a no. less than 4 are completely independent events therefore the conditional probabilies would not hold and they will be the following

P(first die even number | second die <=4) =P(first die even number) = 3/6 =0.5

P(second die <=4|first die even number) =P(second die<=4) =4/6 =0.67